# Training Convolutional Neural Networks with the Forward-Forward algorithm

### 22 Dec 2023 v1

@@ 3.3 Class Activation Maps for explainable AI applications
Class Activation Maps (CAMs) are visual representations that highlight the regions of an input image which provide maximum information to distinguish a specific label. CAMs are obtained by summing the feature maps generated by the convolutional layers, which encode specific spatial characteristics of the image, weighted by their corresponding weights associated with a specific label. These weights quantify how much each considered spatial feature contributes to the recognition of the target class. This technique helps in understanding and interpreting the network behavior in image analysis.
In this context, we exploited each activation value within the hidden convolutional layers used for the classification task. These activations were multiplied by their corresponding weights linked to the image label and obtained from the linear classifier. This architecture enabled the generation of CAMs characterizing the images of the validation dataset. Four examples were reported in Figure 6, obtained by exploiting a 16x16x16 FF-trained CNN architecture, characterized by filters of 5x5 pixels, Adam optimizer, epochs = 200, batch size = 50, learning rate = and varying frequencies (set 1) labeling. The CAMs highlighted which spatial information were processed by the FF-trained CNNs to provide correct classifications. The entire vertical shape was considered to recognize the number 1 (Figure 6 A-B), while three different areas (the upper, the bottom-left and the bottom right parts of the number) were exploited to classify the number 2 (Figure 6 5C-D). The number 7 were recognized basing almost only on its long horizontal line, correlated with the ends of its vertical line (Figures 6 E-F). A digit were classified as a 9 by considering the presence of a round-shape figure, characterized by an appendix in its bottom region.
@@
CAMs also showed that each layer of the FF-trained CNN provides distinct and valuable information for the classification task. This is demonstrated by visualizing the regions independently exploited by each layer of the network to accomplish the classification task (Figure 9, Appendix A). For instance, when considering the digit 7, the second layer of the network provides the most valuable information concerning the inner portion of the horizontal line characterizing the number 7, while the third layer offers more information about the boundaries of the horizontal line.
@@ 4 Discussion and Conclusions
CNNs are considered the gold standard in deep learning-based image analysis. For instance, in biomedical imaging they overcome the drawbacks of subjective analysis in the semi-quantitative visual inspection of samples (Gurcan et al., 2009) and they support experts during their daily clinical routine by reducing their workload (Shmatko et al., 2022). Furthermore, their exploitation of the spatial information within images makes them suitable for the deployment of explainable AI tools (such as Class Activation Maps) which highlight the image regions contributing most significantly to the classification outcome.
Our implementation of FF trained CNN shows that with the right choice of hyperparameters this technique is competitive with backpropagation. These results were obtained without implementing all the possible and suggested optimizations such as enforcing symmetry of the loss function (Lee and Song, 2023) or choosing hard, i.e. easily confused, labels for the negative dataset as suggested by Hinton (2022). We propose that our work shows the potential of FF trained CNNs to address real world computer vision problems. If this technique will in some applications supersede BP is an open question. We believe that this potential exists, especially in the cases of neuromorphic hardware and unsupervised learning.
A better understanding of the FF training will however also expand our understanding of the generic concept of neuronal information processing in all its breadth from biological systems to reservoir computing. The here demonstrated capability of implementing Class Activation Maps provides a first window into these research topics. Achieving deeper insights will also mean to understand how the two innovations of FF, providing positive and negative labels and computing a locally defind goodness parameter, contribute to its success individually and synergetically (Tosato et al., 2023).
@@
This project has received funding from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement No 101034427. The JU receives support from the European Union’s Horizon 2020 research and innovation program and EFPIA. The JU is not participating as a contracting authority in this procurement. This project was also funded by the Ministry for Science and Culture of Lower Saxony as part of the project “Agile, bio-inspired architectures” (ABA).
@@ Appendix A A second way of encoding the labels
For further information, refer to Figure 7.
@@ Appendix B Contribution of the first layer
For further information, refer to Figure 8.
Appendix C Effect of Max Pooling on FF-trained CNNs
We also examined how max pooling layers influence FF-trained CNNs. Specifically, we tested five different configurations. Among these, two exhibited an increase in architectural complexity, characterized by a higher number of filters, which are commonly proposed for MNIST dataset analysis via CNNs. Our analysis, shown in Table 2, indicated a slight decrease in accuracy values when max pooling layers were applied, suggesting that preserving the entire information is preferable. Based on these results, if the conservation of computational resources is a priority, max pooling could be a viable strategy to pursue, especially for architectures with a higher number of filters, as the decrease in CNN performance is limited in such cases. However, for simpler configurations that run more quickly, max pooling significantly affects the measurements, making its application unnecessary.
FF-CNN Network FF Acc. [%] FF + MaxPooling Acc. [%] 128x128x128, filters 7x7 99.14 98.97 16x16x16, filters 5x5 98.16 97.57 64x64x64, filters 3x3 98.30 98.18 32x64x128, filters 7x7 98.90 98.54 16x32x64, filters 7x7 98.47 98.22
Appendix D Additional results CAM
For further information, refer to Figure 9.

### 25 Dec 2023 v2

@@ 3.3 Class Activation Maps
Class Activation Maps (CAMs) are visual representations that highlight the regions of an input image which contribute most to the prediction of a given label. CAMs are obtained by summing up the feature maps generated by the convolutional layers, each weighted with the corresponding weights associated with a specific label. The idea is that each feature map decodes specific spatial characteristics of the input image and the weights quantify how much these characteristics contributes to the recognition of the target class.
In BP trained CNN, CAMs are typically obtained by applying a global average pooling layer after the last convolutional layer and then using the weights connecting this layer to a softmax activation output layer. Here we train a CNN with FF and a linear classifier for inference. The weights of the linear classifier connecting the individual pixels in our feature map with the set of ten output neurons are exactly the weights we need to asses the role of the corresponding pixel for a given prediction.
Figure 6 provides four examples of CAMs of correctly identified images. For the digit 1, (Figure 6 A and B) the entire vertical shape contributes to the correct inference. In contrast, for the digit 2 in Figure 6 C three distinct areas (the upper, the bottom-left and the bottom right parts of the number) contribute to the correct labeling. Similarly expressive regions can be identified for the digits 7 and 9 in Figures 6 E and G.
@@
CAMs also show that the different layers of the FF-trained CNN provide similar but distinct information for the classification task. For instance, when considering the digit 7 in Figure 7, the second layer of the network provides more information on the inner portion of the horizontal line, while the third layer responds more to the boundaries of that horizontal line.
@@ 4 Discussion and Conclusions
CNNs are considered the gold standard in deep learning-based image analysis. For instance, in biomedical imaging, they overcome the drawbacks of subjective analysis in the semi-quantitative visual inspection of samples (Gurcan et al., 2009) and they support experts during their daily clinical routine by reducing their workload (Shmatko et al., 2022). Furthermore, their exploitation of the spatial information within images makes them suitable for the deployment of explainable AI tools (such as Class Activation Maps) which highlight the image regions contributing most significantly to the classification outcome.
Our implementation of FF trained CNN shows that with the right choice of hyperparameters, this technique is competitive with backpropagation. These results were obtained without implementing all the possible and suggested optimizations such as enforcing symmetry of the loss function (Lee and Song, 2023) or choosing hard, i.e. easily confused, labels for the negative data set as suggested by Hinton (2022). We propose that our work shows the potential of FF trained CNNs to address real world computer vision problems. If this technique will in some applications supersede BP is an open question. We believe that this potential exists, especially in the cases of neuromorphic hardware and unsupervised learning.
A better understanding of the FF training will however also expand our understanding of the generic concept of neuronal information processing in all its breadth from biological systems to reservoir computing. The here demonstrated capability of implementing Class Activation Maps provides a first window into these research topics. Achieving deeper insights will also mean to understand how the two innovations of FF, providing positive and negative labels and computing a locally defined goodness parameter, contribute to its success individually and synergetically (Tosato et al., 2023). Moreover, a better understanding why it is beneficial to exclude the first layer during the goodness computation (c.f. Appendix B) would be desirable.
@@
This project has received funding from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement number 101034427. The JU receives support from the European Union’s Horizon 2020 research and innovation program and EFPIA. The JU is not participating as a contracting authority in this procurement. This project was also funded by the Ministry for Science and Culture of Lower Saxony as part of the project “Agile, bio-inspired architectures” (ABA).
@@ Appendix A A Second Way of Encoding the Labels
| FF-CNN Network | Set 1 labels Acc. [%] | Set 2 labels Acc. [%] |
| 128 filters of size 5x5, batch size 15 | 98.98 | 98.95 |
| 64 filters of size 5x5, batch size 15 | 98.83 | 98.81 |
| 32 filters of size 3x3, batch size 50 | 97.95 | 97.94 |
| 32 filters of size 5x5, batch size 50 | 98.50 | 98.47 |
| 32 filters of size 5x5, batch size 100 | 98.40 | 98.40 |
@@ Appendix B Contribution of the First Layer
In order to investigate the effect of the goodness of the first convolutional layer, we train the CNN configurations reported in Section 3.1 again, but this time including the first layer in the goodness computation. Figure 9, reports the results, highlighting that the training of the first layer affects the the speed of convergence of the next layers. Adding the first layer also reduces the overall accuracy of the network by approximately 2%.
Appendix C Effect of Max Pooling
We also examine how max pooling layers influence FF-trained CNNs by testing five different configurations. Two exhibit an increased architectonic complexity by having the number of filters increase for increasing layer number. Our analysis, shown in Table 2, indicates a decrease in accuracy when max pooling layers were applied, suggesting that preserving the entire information is preferable.
| FF-CNN Network | FF Acc. [%] | FF + MaxPooling Acc. [%] |
| 128 filters of size 7x7 | 99.14 | 98.97 |
| 16 filters of size 5x5 | 98.16 | 97.57 |
| 64 filters of size 3x3 | 98.30 | 98.18 |
| 32, 64, 128 filters of size 7x7 | 98.90 | 98.54 |
| 16, 32, 64 filters of size 7x7 | 98.47 | 98.22 |